OpenStudy (anonymous):

differentiate: y= (3x^2 - 4x + 5)^4

OpenStudy (anonymous):

do u know the rules of differentiation

OpenStudy (anonymous):

do you?

OpenStudy (anonymous):

u must write w.r.t. to what

OpenStudy (anonymous):

Let \[\Large t=3x^2-4x+5\] \[\Large \implies y=t^4\] now differentiate y with respect to t and differentiate t with respect to x , then use chain rule \[\Large \frac{dy}{dx}=\frac{dy}{dt}*\frac{dt}{dx}\] \[\Large \frac{dy}{dt}=4t^3\] \[\Large \frac{dt}{dx}=6x-4\] just use chain rule now \[\Large \frac{dy}{dx}=\frac{dy}{dt}*\frac{dt}{dx}\]

OpenStudy (anonymous):

y=f(g(x)) dy/dx = f'(x) x g'(x) so y=(3x^2-4x+5)^4 dy/dx= 4(3x^2-4x+5) (6x-4)

OpenStudy (anonymous):

hmm.. what is the final answer?

OpenStudy (anonymous):

dy/dx= 4(3x^2-4x+5) (6x-4) in factorised form

OpenStudy (anonymous):

4(3x^2-4x+5)^3 (6x-4) opps sorry missed a ^3

OpenStudy (anonymous):

then after that, i will simplify that?

OpenStudy (anonymous):

usually you don't need to~ they usually prefer answer in factorised form~ so that will be the final answer :)

OpenStudy (anonymous):

ahh! thank you so much .. now, i know!